OpenStudy (anonymous):

Solve the exponential equation. 9^2x = 27

OpenStudy (anonymous):

\[9=3^2,27=3^3\] so you can rewrite this as \[3^{4x}=3^3\] and then solve \(4x=3\)

OpenStudy (anonymous):

so would the answer be .75